An aqueous solution has 18% of solute. By chemical analysis, it was known that it has 10 mol% solute.
What is the molecular weight of the solute?
1
Expert's answer
2020-10-21T12:51:52-0400

m = n*M, m - weight, n - quantity of substance, M - molecular weight

mx /18 = mH2O / 82

nx = nH2O / 9

nx *Mx = nH2O * MH2O*18/82

Mx = MH2O*18*9/82 = 18*18*9/82=35.56